# e-Profits: A Business-Aligned Evaluation Metric for Profit-Sensitive Customer Churn Prediction

### v1

@@ 4.2 Maven Dataset Analysis
We now turn to the Maven Telecom dataset, which features a richer set of features and a slightly different class distribution compared to IBM. We estimate customer retention probabilities using Kaplan-Meier analysis. The average retention rate in the Maven dataset is approximately 75%. Figure 4 summarises the retention estimates, lift behaviour, and classification performance.
(a) Kaplan-Meier estimated retention rate
@@
The top-decile lift reaches 3.46, indicating effective risk segmentation. The ROC curve confirms that EBM achieves the highest AUC again, which is consistent with the findings of the IBM dataset.
@@ 4.2.1 Classifier Performance across Metrics
Classifier performance on the Maven dataset mirrors many trends observed on the IBM dataset. Full results are shown in Tables 9 to 13, comparing five tuning strategies and evaluating both global and top-segment e-Profits under ARR and TRR.
As with IBM, AUC-optimised models (Table 9) rank well in traditional metrics, but LightGBM and XGBoost surpass EBM in top-segment e-Profits. EMP-tuned models (Table 12) again fail to generalise well in high-value scenarios.
Most notably, Table 13 and Figure 5 demonstrates that when tuned directly on e-Profits, EBM yields the highest total profitability, while LightGBM captures the highest returns in the top segment. This reinforces e-Profits as a practical, fine-grained tool for selecting models that maximise business value.
Table 9: Classifier performance on Maven dataset when tuned for AUC: Includes traditional metrics, lift indicators, and business-aligned evaluation using e-Profits (ARR/TRR) and EMP under both full and top-segment scenarios.
Name F1 (%) Acc. (%) AUC (%) EMP TDL Lift Index e-Prof (0.2, 0.3, ARR) e-Prof (0.2, 0.3, TRR) e-Prof (1.0, 0.3, ARR) e-Prof (1.0, 0.3, TRR) EMP (0.2, 200, 10, 1) EMP (1.0, 200, 10, 1)
RF 73.71 86.65 80.18 13.09 3.4332 0.8521 6454 7533 28624 64093 47.18 13.09
GBC 75.00 86.75 81.45 13.25 3.4508 0.8589 7608 8272 30449 64715 48.084 13.25
CBC 75.19 86.95 81.49 13.44 3.4500 0.8616 7173 7910 30039 64919 48.03 13.44
LGBM 75.88 87.10 82.12 13.37 3.4500 0.8600 7874 8646 30971 65744 47.18 13.38
XGB 76.14 87.25 82.28 13.28 3.4700 0.8588 7704 8246 31370 66169 47.38 13.28
EBM 76.40 87.41 82.44 13.47 3.4684 0.8620 8009 8602 31412 66514 48.54 13.47
Table 10: Classifier performance on Maven dataset when tuned for Accuracy: Highlights classification quality alongside financial impact measured by e-Profits and EMP across customer segments.
Name F1 (%) Acc. (%) AUC (%) EMP TDL Lift Index e-Prof (0.2, 0.3, ARR) e-Prof (0.2, 0.3, TRR) e-Prof (1.0, 0.3, ARR) e-Prof (1.0, 0.3, TRR) EMP (0.2, 200, 10, 1) EMP (1.0, 200, 10, 1)
RF 73.26 86.60 79.78 13.14 3.4156 0.8532 6619 7636 27969 62905 47.35 13.14
GBC 75.00 86.75 81.45 13.25 3.4508 0.8589 7608 8272 30449 64715 48.08 13.25
CBC 75.19 86.95 81.49 13.44 3.4508 0.8616 7173 7910 30039 64919 48.03 13.44
XGB 75.41 87.00 81.68 13.41 3.4684 0.8626 7125 7636 30495 64987 48.54 13.41
LGBM 75.88 87.10 82.12 13.37 3.4508 0.8598 7874 8646 30971 65744 47.18 13.37
EBM 76.40 87.41 82.44 13.47 3.4683 0.8619 8009 8602 31412 66514 48.54 13.47
Table 11: Classifier performance on Maven dataset when tuned for F1-score: Shows how financial metrics like e-Profits shift model preferences compared to standard evaluation.
Name F1 (%) Acc. (%) AUC (%) EMP TDL Lift Index e-Prof (0.2, 0.3, ARR) e-Prof (0.2, 0.3, TRR) e-Prof (1.0, 0.3, ARR) e-Prof (1.0, 0.3, TRR) EMP (0.2, 200, 10, 1) EMP (1.0, 200, 10, 1)
RF 73.71 86.65 80.18 13.09 3.4332 0.8521 64544 7533 28624 64093 47.18 13.09
GBC 75.00 86.75 81.45 13.25 3.4508 0.8589 7608 8272 30449 64715 48.08 13.25
CBC 75.19 86.95 81.49 13.44 3.4508 0.8616 7173 7910 30039 64919 48.03 13.44
XGB 75.86 87.25 81.97 13.37 3.4684 0.8604 7366 7885 30387 64963 48.25 13.37
LGBM 75.88 87.10 82.12 13.37 3.4508 0.8598 7909 8646 30971 65744 47.18 13.37
EBM 76.40 87.41 82.44 13.47 3.4684 0.8619 8008 8602 31412 66514 48.54 13.47
Table 12: Classifier performance on Maven dataset when tuned for EMP: EMP’s reliance on fixed assumptions is contrasted with e-Profits’ granular evaluation across ARR and TRR.
Name F1 (%) Acc. (%) AUC (%) EMP TDL Lift Index e-Prof (0.2, 0.3, ARR) e-Prof (0.2, 0.3, TRR) e-Prof (1.0, 0.3, ARR) e-Prof (1.0, 0.3, TRR) EMP (0.2, 200, 10, 1) EMP (1.0, 200, 10, 1)
XGB 47.94 80.22 65.75 12.79 3.4508 0.8453 0.00 0.00 9311 34366 44.97 12.79
GBC 58.48 82.90 70.66 12.77 3.4156 0.8460 1974 2145 15567 43252 46.16 12.77
CBC 66.59 84.37 75.35 12.76 3.3980 0.8475 3262 3763 218045 54852 45.48 12.76
RF 73.41 86.44 80.04 12.91 3.4156 0.8518 6567 7601 28318 63747 47.18 12.91
LGBM 75.88 87.10 82.12 13.37 3.4508 0.8598 7909 8646 30971 65744.38 47.18 13.37
EBM 76.40 87.41 82.44 13.47 3.4684 0.8619 8008 8602 31412 66514 48.54 13.47
Table 13: Classifier performance on Maven dataset when tuned directly for e-Profits: Demonstrates effectiveness of personalised financial evaluation using e-Profits under different retention assumptions.
Name F1 (%) Acc. (%) AUC (%) EMP TDL Lift Index e-Prof (0.2, 0.3, ARR) e-Prof (0.2, 0.3, TRR) e-Prof (1.0, 0.3, ARR) e-Prof (1.0, 0.3, TRR) EMP (0.2, 200, 10, 1) EMP (1.0, 200, 10, 1)
RF 73.71 86.65 80.18 13.09 3.4332 0.8521 6454 7533 28624 64093 47.18 13.09
GBC 75.00 86.75 81.45 13.25 3.4508 0.8589 7608 8272 30449 64715 48.08 13.25
CBC 75.19 86.95 81.49 13.44 3.4508 0.8616 7173 7910 30040 64919 48.03 13.44
LGBM 75.88 87.10 82.12 13.37 3.4508 0.8598 7909 8646 30971 65744 47.18 13.37
XGB 76.14 87.25 82.28 13.28 3.4684 0.8588 7704 8246 31370 66169 47.38 13.28
EBM 76.40 87.41 82.44 13.47 3.4684 0.8619 8008 8602 31412 66514 48.54 13.47
4.2.2 Key Insights
  • Results from the Maven dataset reinforce the limitations of traditional metrics for model selection in profit-sensitive scenarios.
  • EMP tuning, although profit-aware, is constrained by its fixed assumptions and fails to highlight top-performing models in high-value segments.
  • e-Profits again provides a clearer differentiation, especially for top-segment profitability, helping identify models like LightGBM that may otherwise be overlooked.
  • The consistency of findings across datasets highlights e-Profits as a reliable, business-aligned evaluation framework.
Figure 5: Radar graph for Maven dataset (Tuned on e-profits considering 20% profitable customers and tenure based retention)
5 Discussion
@@
Our findings across the IBM and Maven datasets reveal consistent patterns that underscore the practical value of using e-Profits as an evaluation metric for churn prediction models. Traditional performance metrics such as AUC and F1-score did not consistently identify the most financially effective models. For instance, classifiers like LGBM and CatBoost, which did not top the rankings under accuracy or AUC, emerged as highly profitable when evaluated using e-Profits. This reordering highlights a critical discrepancy between predictive correctness and financial value, underscoring the limitations of relying solely on statistical metrics for real-world business decisions.
@@
Moreover, e-Profits proved especially effective in identifying models that excel within high-value customer segments. In both datasets, CatBoost and GBC generated the highest profits when analysis was focused on the top 20% of customers sorted by predicted churn risk and estimated profitability. These models were not necessarily the best according to traditional metrics, reinforcing the importance of adopting evaluation criteria that prioritise return on investment (ROI) over aggregate accuracy. The ability to tailor model evaluation towards strategic, high-impact interventions makes e-Profits a valuable tool for resource-constrained retention campaigns.
@@
A key insight from our analysis is the limited utility of commonly used lift metrics in reflecting actual business outcomes. While lift remains a popular heuristic in industry, it often fails to account for the heterogeneity in customer value and the variable cost of retention. For example, Table 7 illustrates scenarios where models with strong lift scores delivered zero e-Profits, indicating that churn likelihood alone is an insufficient criterion for profitable targeting. By combining churn risk with revenue potential and cost structure, e-Profits offers a more holistic and financially meaningful basis for model selection.
In addition to these strategic insights, we also note several practical considerations in the implementation of e-Profits. First, accurate estimation of customer lifetime value and retention probabilities requires access to detailed data, including monthly revenue, tenure, and timestamped churn labels. Organisations may need to enrich existing customer datasets to support such granular modelling. Second, e-Profits is sensitive to cost parameters such as contact cost and the cost of providing retention offers. While our experiments employed domain-informed heuristics for offer costs (e.g., proportions of CLV informed by CRM practices), businesses may need to calibrate these parameters based on internal financial policies and strategic goals. Third, our current approach assumes a binary intervention model, deciding whether or not to contact a customer based on predicted churn. Future extensions could consider multi-tiered or continuous intervention strategies to capture more complex business realities.
We also justified the use of Kaplan-Meier survival analysis for retention rate estimation, citing its simplicity, non-parametric nature, and understandability. These qualities make it suitable for business contexts where model transparency and ease of implementation are key. Although more complex methods such as DeepSurv [26] and DRSA [38] may offer improved predictive accuracy, their integration typically requires changes to model training workflows and increased computational overhead. For the purposes of post hoc model evaluation, Kaplan-Meier offers a robust and practical trade-off, delivering reliable customer-level retention estimates without excessive complexity.
In summary, e-Profits extends beyond predictive performance to deliver insights directly aligned with financial decision-making. Its application enables marketing and analytics teams to select models that maximise business impact rather than statistical precision alone, positioning it as a valuable addition to the model evaluation toolkit in customer retention contexts.

### v2

@@ 4.2 Maven Dataset Analysis
We now turn to the Maven Telecom dataset, which features a richer set of customer attributes and a slightly different class distribution compared to IBM. Customer retention probabilities are estimated using Kaplan-Meier survival analysis. The ARR estimated from the training split of the Maven dataset is approximately 0.75. As with IBM, the Kaplan-Meier curve is fitted only on the training data and then applied unchanged during test-set evaluation. Figure 3 summarises the retention estimates, lift behaviour, and classification performance. The top-decile lift reaches 3.46, indicating effective risk segmentation. Consistent with the IBM dataset, the ROC curve shows that EBM attains the highest AUC.
(a) Kaplan-Meier estimated retention curve
@@ 4.2.1 Classifier Performance across Metrics
Classifier performance on the Maven dataset mirrors many of the trends observed for the IBM dataset. Full results are reported in Table 4, comparing five tuning strategies and evaluating both full-population and top-segment e-Profits under ARR and TRR.
Models optimised for AUC perform strongly on traditional predictive metrics; however, Random Forest achieves the highest e-Profits under AUC tuning, indicating that strong ranking performance does not necessarily translate into maximum profitability. When models are tuned directly on e-Profits (Figure 5), XGB delivers the highest total (full-population) profitability, while LightGBM captures the highest returns in the top 20% customer segment.
More generally, XGB achieves the highest full-population e-Profits when tuned on Accuracy, F1-score, and e-Profits, and ranks second under EMP tuning, where EBM attains a slightly higher profit. In contrast, LightGBM consistently outperforms other classifiers on top-segment e-Profits when tuned on F1-score, EMP, and e-Profits, and ranks second when optimised for Accuracy. Consistent with the IBM dataset, directly optimising models for e-Profits yields the highest overall profitability in the Maven dataset while preserving strong AUC and F1-score performance.
The radar plots shown by Figures 4 and 5 ((Higher values indicate better performance) further illustrate that no single classifier dominates across both predictive and profit-based metrics. Together, these results reinforce e-Profits as a practical and fine-grained evaluation criterion for selecting models that maximise business value under different targeting objectives.
Table 4: Classifiers’ performance on the Maven dataset under AUC, Accuracy, F1-score, EMP e-Profits, respectively.
Name F1 (%) Acc. (%) AUC (%) TDL Lift Index e-Prof (0.2, 0.3, ARR) e-Prof (0.2, 0.3, TRR) e-Prof (1.0, 0.3, ARR) e-Prof (1.0, 0.3, TRR) EMP (0.2, 200, 10, 1) EMP (1.0, 200, 10, 1)
Tuned on AUC
RF 73.41 86.44 92.66 3.4156 0.8518 271168 286715 969408 996582 47.18 12.91
GBC 68.48 82.90 91.87 3.4156 0.8460 67189 80966 530295 381167 46.16 12.77
XGB 73.94 80.22 92.07 3.4508 0.8470 0 0 320168 140032 45.14 12.82
CBC 72.59 84.37 92.21 3.3980 0.8475 111043 137715 742668 596238 45.48 12.76
LGBM 70.20 85.79 93.56 3.4684 0.8582 171384 213167 835585 726076 47.52 13.31
EBM 76.11 87.20 94.16 3.4684 0.8607 223571 239644 969120 880083 47.72 13.42
Tuned on Accuracy
RF 73.26 86.60 92.99 3.4156 0.8532 225359 289009 952522 868474 47.35 13.14
GBC 75.00 86.75 93.78 3.4508 0.8589 259017 321128 1036960 966037 48.08 13.25
XGB 75.41 87.00 94.12 3.4684 0.8612 268458 337911 1034441 958907 48.37 13.33
CBC 75.19 86.95 94.31 3.4508 0.8616 244202 309784 1023033 948466 48.03 13.44
LGBM 73.74 86.60 94.07 3.4684 0.8611 217283 276479 967808 883372 48.03 13.44
EBM 76.33 87.35 94.30 3.4683 0.8626 250647 316490 1072989 1003393 47.69 13.50
Tuned on F1-score
RF 73.71 86.65 92.84 3.4332 0.8521 219722 279940 974823 887795 47.18 13.09
GBC 75.00 86.75 93.78 3.4508 0.8589 259017 321128 1036960 966037 48.08 13.25
XGB 75.41 87.10 94.12 3.4684 0.8612 250647 316490 1034441 958907 48.37 13.33
CBC 76.19 86.95 94.31 3.4508 0.8616 244202 309784 1023033 948466 48.03 13.44
LGBM 75.88 87.10 93.84 3.4508 0.8598 268075 342257 1094763 996903 47.18 13.37
EBM 76.19 87.10 94.12 3.4508 0.8607 265768 333044 1075533 944924 48.25 13.42
Tuned on EMP
RF 73.51 86.55 92.87 3.4332 0.8526 217001 278956 955944 867898 47.26 13.12
GBC 74.83 86.60 93.67 3.4508 0.8596 232111 296995 1026399 956475 47.35 13.29
XGB 75.74 87.20 94.21 3.4684 0.8623 263369 335457 1045890 980577 47.86 13.39
CBC 76.22 87.41 94.15 3.4508 0.8611 249991 315433 1056558 982833 48.54 13.39
LGBM 75.88 87.10 93.84 3.4508 0.8598 268075 342257 1054763 996903 47.18 13.37
EBM 75.50 86.90 94.08 3.4684 0.8612 258579 319910 1052440 971231 47.52 13.43
Tuned on e-Profits (under TRR on full population)
RF 73.71 86.65 92.84 3.4332 0.8521 219722 279940 994823 987795 47.18 13.09
GBC 75.00 86.75 93.78 3.4508 0.8589 259017 321128 1066960 1046037 48.08 13.25
XGB 76.30 87.30 93.70 3.4684 0.8600 231403 294114 1092216 1063393 48.54 13.28
CBC 75.19 86.95 94.31 3.4508 0.8616 244202 309784 1053033 1048466 48.03 13.44
LGBM 75.80 87.05 93.73 3.4684 0.8582 267352 340635 1080897 1025717 48.03 13.27
EBM 76.19 87.10 94.12 3.4508 0.8607 265768 333044 1075533 1004924 48.25 13.43
e-Profits are reported as $e\text{-}Prof(\eta,M,\mathrm{RR})$, where $\eta\in\{0.2,1.0\}$ denotes the evaluated customer fraction (top 20% or full population), $M=0.3$ is the profit margin used in Eq. 3, and $\mathrm{RR}\in\{\mathrm{ARR},\mathrm{TRR}\}$ specifies whether average or tenure-based retention probabilities are used. EMP values are computed as $\mathrm{EMP}(\eta,\mathrm{CLV},C_{\text{offer}},C_{\text{contact}})$ with $\mathrm{CLV}=200$, $C_{\text{offer}}=10$, and $C_{\text{contact}}=1$.
*CBC=CatBoost, Acc.=Accuracy, TDL=Top-Decile Lift
Figure 4: Radar graph for the IBM dataset optimised for e-Profits on full population under TRR (axes show normalised scales for AUC, F1, Accuracy, EMP, and e-Profits)
Figure 5: Radar graph for Maven dataset optimised for e-Profits on full population under TRR (axes show normalised scales for AUC, F1, Accuracy, EMP, and e-Profits).
4.3 Statistical Significance, Uncertainty, and Model Re-Ranking
To quantify uncertainty and assess whether observed differences are statistically meaningful, we report bootstrap confidence intervals and conduct paired non-parametric statistical tests. Our analysis focuses on e-Profits, and all models are evaluated on held-out test sets for both the Maven and IBM datasets.
Table 5 reports test-set performance for the best-performing models selected under different tuning objectives. Predictive metrics such as AUC, F1-score, and Accuracy remain broadly comparable across tuning policies, with differences typically within one to two percentage points. In contrast, e-Profits varies substantially depending on the tuning objective. Importantly, selecting models using e-Profits does not degrade predictive performance on either dataset, while consistently yielding the highest e-Profits. This demonstrates that conventional classification metrics are insufficient proxies for downstream business value and that profit-aware evaluation provides complementary and decision-relevant information.
To assess estimation variability, we perform bootstrap resampling of the test sets (500 iterations) and report bootstrap means and 95% confidence intervals for expected profit in Table 6. Predictive metrics exhibit relatively narrow confidence intervals, indicating stable estimation, whereas profit-based metrics display wider intervals, reflecting the inherently higher variability of economic outcomes. Despite this variability, models selected under e-Profits consistently achieve the highest mean profit with confidence intervals that dominate those of models selected using traditional objectives, indicating that the observed profit gains are not attributable to sampling noise.
4.3.1 Statistical Significance of Profit Differences
To formally test whether differences in expected profit are statistically significant, we conduct paired Wilcoxon signed-rank tests across identical model families evaluated on the same test sets. These tests operate on paired profit differences and assess whether profit gains under e-Profits are systematic rather than random. Results are summarized in Table 7. Across both the Maven and IBM datasets, models selected using e-Profits significantly outperform those tuned for AUC and Accuracy in terms of full-population e-Profits ($p\ll 0.001$). The magnitude of these gains is substantial on both datasets, confirming that profit improvements under e-Profits are statistically significant and consistent across model families.
4.3.2 Model Re-Ranking Under Profit-Based Evaluation
While predictive metrics often rank models similarly, evaluating models using expected profit induces systematic reordering. To quantify this effect, we compute Spearman’s rank correlation between AUC-based rankings and profit-based rankings. Statistical significance is assessed via paired Wilcoxon signed-rank tests on rank differences.
Results are reported in Table 8 showing agreement between AUC-based and profit-based rankings is moderate rather than perfect on both the datasets. On the Maven dataset, Spearman’s rank correlation is approximately $\rho=0.49$ for full-population e-Profits and $\rho=0.45$ for top-20% targeting. On IBM, rank agreement is slightly lower but comparable, with $\rho\approx 0.43$ for full-population profit and $\rho\approx 0.39$ for top-20% profit. In all cases, paired Wilcoxon signed-rank tests on rank differences strongly reject perfect concordance ($p<10^{-23}$ for Maven and $p<10^{-18}$ for IBM), confirming that profit-based evaluation induces systematic and statistically significant re-ranking rather than random variation. These results indicate that predictive performance metrics capture only part of the information relevant for economic decision-making, while profit-aware evaluation provides complementary guidance, particularly for targeted retention strategies.
Taken together, these results demonstrate that business-aware evaluation using e-Profits leads to statistically justified model selection decisions and materially different rankings than traditional predictive metrics. These effects are consistent across datasets with differing characteristics, reinforcing the importance of profit-aware evaluation in applied decision-making contexts.
Table 5: Test-set performance of best-performing models selected under different tuning objectives for the Maven and IBM datasets.
Dataset Policy Model AUC (%) F1 (%) Acc. (%) e-Profits (Full TRR)
Maven AUC RF 92.63 73.41 86.44 996582
Accuracy EBM 94.30 76.33 87.35 1003393
F1 LGBM 93.84 75.88 87.10 996903
EMP LGBM 93.84 75.88 87.10 996903
e-Profits XGB 93.70 76.30 87.30 1063393
IBM AUC RF 83.77 57.20 79.18 9104621
Accuracy XGB 85.48 59.32 80.27 9953379
F1 XGB 85.58 59.51 80.36 9953379
EMP LGBM 84.68 56.87 79.60 9829970
e-Profits XGB 83.82 58.47 79.22 10991460
*For each tuning policy, the model column reports the classifier that achieves the highest e-Profits on the held-out test set among all models trained under that policy. The reported AUC, F1-score, and Accuracy values correspond to this profit-maximising model and are not necessarily the maximum values achievable under the tuning objective.
Table 6: Bootstrap mean and 95% confidence intervals (500 resamples) for e-Profits (Full TRR) on Maven and IBM dataset.
Dataset Policy Mean CI Low CI High
Maven AUC 996925 862389 1127743
Accuracy 1003816 871303 1137394
e-Profits 1061128 875735 1129249
IBM AUC 9103018 8109360 10180288
Accuracy 9952973 8923399 11018421
e-Profits 10913261 8945986 11217734
Table 7: Paired Wilcoxon signed-rank tests comparing bootstrap mean differences in full-population TRR-based e-Profits across identical model families under different tuning objectives.
Dataset Policy A Policy B Metric Mean Diff $p$-value
Maven e-Profits AUC e-Profits (Full TRR) +64203 $\ll 0.001$
e-Profits Accuracy e-Profits (Full TRR) +57312 $\ll 0.001$
IBM e-Profits AUC e-Profits (Full TRR) +1810243 $\ll 0.001$
e-Profits Accuracy e-Profits (Full TRR) +960288 $\ll 0.001$
Table 8: Agreement between AUC-based ranking and profit-based ranking, measured by Spearman’s $\rho$.
Dataset Profit metric Spearman $\rho$ (95% CI) $p$-value
Maven eProfits (Full TRR) $\approx 0.49$ $10^{-24}$
eProfits (Top20 TRR) $\approx 0.45$ $10^{-23}$
IBM eProfits (Full TRR) $\approx 0.43$ $1.50\times 10^{-18}$
eProfits (Top20 TRR) $\approx 0.39$ $7.07\times 10^{-19}$
5 Discussion and Implications
@@
Our findings across the IBM and Maven datasets reveal consistent patterns that underscore the practical value of using e-Profits as an evaluation metric for churn prediction models. Traditional performance metrics such as AUC and F1-score do not consistently identify the most financially effective models. For instance, classifiers such as LightGBM and CatBoost, which do not always top the rankings under Accuracy or AUC, emerge as highly profitable when evaluated using e-Profits. This reordering highlights a critical discrepancy between predictive correctness and financial value, underscoring the limitations of relying solely on statistical metrics for real-world business decisions.
@@
Moreover, e-Profits proves especially effective in identifying models that excel within high-value customer segments. Across both datasets, models such as LightGBM and XGBoost frequently deliver the highest profitability when evaluation is restricted to the top 20% of customers ranked by expected profit. These models are not necessarily preferred under traditional metrics, reinforcing the importance of evaluation criteria that prioritise return on investment (ROI) over aggregate predictive accuracy. The ability to tailor model evaluation toward strategic, high-impact interventions makes e-Profits particularly suitable for resource-constrained retention campaigns.
These results confirm that explicitly coupling churn risk with heterogeneous customer value and intervention cost leads to systematically different model rankings than accuracy-based evaluation. Models with similar AUC or F1-score can differ substantially in how they prioritise high-CLV customers or avoid costly false positives. By translating predictions into per-customer economic outcomes, e-Profits amplifies these differences and reveals profit-relevant distinctions that remain invisible to purely statistical metrics.
@@
A further insight from our analysis concerns the limited utility of commonly used lift-based measures in reflecting actual business outcomes. While lift remains a popular heuristic in industry, it does not account for heterogeneity in customer value or retention cost. For example, Table 3 illustrates scenarios in which models with strong lift scores nevertheless deliver zero e-Profits, indicating that churn likelihood alone is an insufficient criterion for profitable targeting. By jointly considering churn risk, revenue potential, and intervention cost, e-Profits provides a more holistic and economically meaningful basis for model selection.
From an implementation perspective, e-Profits requires reliable estimation of customer lifetime value and retention probabilities, which in turn depends on the availability of revenue, tenure, and churn-timing data. In addition, profitability outcomes are sensitive to cost parameters such as contact and offer costs, which must be calibrated to organisational policies and campaign constraints. While our experiments adopt domain-informed heuristics, these parameters can be readily adapted to reflect business-specific assumptions. Our current formulation assumes a binary intervention decision; extending the framework to support multi-tier or continuous intervention strategies represents a natural extension.
Our results do not imply that traditional metrics such as AUC or F1-score are obsolete. These metrics remain appropriate when the primary objective is ranking quality or classification accuracy under uniform cost assumptions. However, when customer value is heterogeneous, intervention costs scale with CLV, or budgets are constrained, profit-aware evaluation becomes essential. In such settings, e-Profits offers a more decision-relevant criterion for model selection.
By explicitly aligning model evaluation with customer value, retention probability, and intervention cost, e-Profits provides a decision-oriented criterion that better reflects operational objectives. While profit-based alternatives such as EMP move in this direction, their reliance on fixed population-level assumptions limits their effectiveness in heterogeneous customer bases.
5.1 Managerial Implications, Applicability
From a managerial perspective, our findings indicate that selecting churn models solely based on predictive metrics such as AUC or F1-score can lead to systematically suboptimal retention strategies, particularly under budget constraints. In practice, organisations rarely intervene on all predicted churners; instead, retention campaigns must allocate limited resources across customers with heterogeneous value. As a result, model evaluation criteria that ignore customer value and intervention costs can misguide operational decision-making.
Deploying models evaluated with e-Profits enables firms to (i) select different models for mass targeting versus high-value, budget-limited campaigns, (ii) prioritise models that reduce costly false positives among low-value customers, and (iii) explicitly align model evaluation with campaign economics rather than abstract predictive performance. This allows data science teams to justify model deployment decisions in financial terms that are directly interpretable and actionable for marketing, CRM, and customer success stakeholders. Although absolute profit values naturally scale with the choice of cost parameters, we empirically observed that relative model rankings remained stable across reasonable parameter ranges (e.g., $\mathrm{CPO}\in[0.05,0.2]$). This observation motivated our decision to focus on ranking robustness rather than exhaustive sensitivity grids in the present study.
Although our empirical evaluation uses telecom datasets, the e-Profits framework is not tied to telecom-specific features. In practice, customer lifetime value and retention can be derived from basic billing and tenure information (e.g., monthly revenue and time-to-churn or subscription length), while intervention costs are specified through business assumptions rather than inferred directly from the data. This makes e-Profits applicable to other subscription-based domains such as banking or SaaS, provided that comparable revenue and tenure signals are available and an operational definition of churn exists.
The main challenges are therefore practical rather than methodological. In many non-telecom settings, suitable customer-level data are not publicly available for experimentation, and churn dynamics (e.g., soft churn or account dormancy) may require domain-specific calibration of value and cost components. Nevertheless, these challenges concern data availability and business calibration rather than limitations of the evaluation framework itself.
While our experiments use fixed cost and margin values, e-Profits is inherently parametric. Varying these parameters is expected to rescale profits and shift targeting thresholds without altering the underlying decision logic, allowing organisations to calibrate the metric to their own cost structures.
5.2 Limitations and Future Directions
This study positions e-Profits as a decision-oriented evaluation and model-selection metric rather than a causal treatment-effect estimator. We assume that intervening on a true churner successfully prevents churn and do not explicitly model heterogeneous responsiveness to retention actions. As a result, e-Profits evaluates the profitability of a targeting policy under explicit cost and retention assumptions rather than estimating incremental profit relative to a no-intervention counterfactual. Integrating causal uplift or treatment-effect modelling to estimate customer-specific intervention effectiveness therefore represents an important direction for future work.
A further limitation concerns the estimation of retention probabilities using Kaplan-Meier survival analysis. While Kaplan-Meier is non-parametric, transparent, and computationally lightweight, it assumes non-informative censoring and captures retention heterogeneity only through customer tenure using a single population-level survival curve. In settings where churn dynamics depend strongly on customer covariates or evolve over time, this assumption may be restrictive. Future work may therefore explore more expressive survival models to capture heterogeneity beyond tenure, including parametric approaches (e.g., Weibull) and covariate-aware models such as Cox proportional hazards with time-varying covariates. Marketing-oriented parametric churn models such as Beta-Geometric / BG-NBD formulations also represent a promising alternative, offering smoother and more stable retention estimates in sparse or long-tenure regions.
Finally, while our experiments focus on telecom datasets, the e-Profits framework is not domain-specific. Applying the metric in other subscription-based contexts (e.g., banking or SaaS) would require domain-calibrated definitions of churn, revenue, and intervention costs. Investigating domain transfer, parameter calibration, and extensions to multi-tier or continuous intervention strategies represents an important avenue for future research.